### v2

# CVOCSemRPL: Class-Variance Optimized Clustering, Semantic Information Injection and Restricted Pseudo Labeling based Improved Semi-Supervised Few-Shot Learning

Abstract
Few-shot learning has been extensively explored to address problems where the amount of labeled samples is very limited for some classes. In the semi-supervised few-shot learning setting, substantial quantities of unlabeled samples are available. Such unlabeled samples are generally cheaper to obtain and can be used to improve the few-shot learning performance of the model. Some of the recent methods for this setting rely on clustering to generate pseudo-labels for the unlabeled samples. Since the effectiveness of clustering heavily influences the labeling of the unlabeled samples, it can significantly affect the few-shot learning performance. In this paper, we focus on improving the representation learned by the model in order to improve the clustering and, consequently, the model performance. We propose an approach for semi-supervised few-shot learning that performs a class-variance optimized clustering coupled with a cluster separation tuner in order to improve the effectiveness of clustering the labeled and unlabeled samples in this setting. It also optimizes the clustering-based pseudo-labeling process using a restricted pseudo-labeling approach and performs semantic information injection in order to improve the semi-supervised few-shot learning performance of the model. We experimentally demonstrate that our proposed approach significantly outperforms recent state-of-the-art methods on the benchmark datasets.
1 Introduction
Modern deep networks often surpass human performance but demand ever deeper architectures and vast labeled datasets, whose annotation is costly and time-consuming. This data bottleneck motivates few-shot learning, where models must generalize from only a handful of examples. Therefore, researchers have been looking at few-shot learning (FSL) as a solution. Few-shot learning tackles new classes with only a handful of labeled examples by first training a feature extractor on richly annotated base classes and then transferring those learned representations to extract meaningful embeddings for the scarce few-shot samples. The performance of existing FSL methods is still limited due to the extremely limited availability of labeled samples [1, 2]. Unlabeled data is often plentiful and easy to collect, so incorporating large pools of unlabeled examples alongside the scarce few-shot samples, known as semi-supervised few-shot learning, can substantially boost performance. In this paper, we tackle the semi-supervised few-shot learning (SSFL) problem.
Researchers have proposed multiple approaches to address the SSFSL problem, such as clustering to assign pseudo-labels to unlabeled samples and utilizing them to improve the FSL performance [3, 4]. Clustering relies heavily on the quality of representation learned by the network, especially on the base classes, since this ensures that newly observed images, even from unseen classes, are far away from images of other classes and close to images from the same class in the representation space. A model with better quality representation will facilitate better clustering of extracted features and consequently lead to better pseudo-label assignment to the unlabeled samples (refer to Table 5).
We propose a novel SSFSL approach in this work. Specifically, we use a class variance optimized clustering (CVOC) process to cluster given images and use a semantic injection network to refine the cluster centers by incorporating semantic information. The proposed CVOC process optimizes intra- and inter-class separation and utilizes a cluster separation tuner to achieve improved clustering (refer to Table 5). We use the refined cluster centers to perform restricted pseudo-labeling of the unlabeled samples and incorporate all labeled samples to make predictions for the query samples (see Sec. 4). We perform experiments on multiple benchmark datasets and demonstrate that our proposed approach significantly outperforms the compared approaches.
The key contributions of this paper are as follows:
  1. 1.
    We propose a novel semi-supervised few-shot learning approach that outperforms the recent methods in this setting, as shown in Tables 123.
  2. 2.
    We experimentally demonstrate how our approach improves the pseudo-labeling accuracy, as shown in Table 5, and is, therefore, able to outperform the compared methods.
  3. 3.
    We experimentally demonstrate how incorporating semantic features of classes and managing the intra-class compactness and inter-class separation helps to further improve the performance of the model, as shown in Table 5.
2 Related Works
2.1 Few-shot Image Classification
Few-shot learning (FSL) methodologies [5, 6, 7, 8, 9] generally pre-train the model using extensive data from the base classes and subsequently modify the pre-trained model for recognizing novel categories.  [10] utilizes an average feature prototype of the classes with very few samples and performs predictions by identifying the nearest prototype. TADAM [11] adapts the features of the few-shot class images using the few support images available for such classes. [12] improves upon prototypical network using an importance-weighting strategy for the different support samples of a class and makes the prototypes more informative. Meta-learning methods learn to quickly adapt to new classes using very few samples, and these methods involve training a model at a specific level as well as at an abstract level, referred to as learning to learn.  [13]. Researchers have proposed various meta-learning-based FSL approaches, such as  [14, 15].
2.2 Pseudo-labeling based Semi-supervised Few-shot Methods
The semi-supervised few-shot learning (FSL) setting is a modified FSL setting with access to unlabeled samples. Several semi-supervised FSL methods assign pseudo-labels to unlabeled samples and then consider these samples as labeled samples for the few-shot classification task [16, 17, 13]. However, if some of the pseudo labels are incorrect, the performance of the model is degraded. Some methods utilize label propagation to generate pseudo-labels [18][19], selectively sample high-confidence labels, but treat each sample independently and miss out on broader relational context.  [20] attempts to unify pseudo-labeled data in a single metric space, but also does not consider the inter-sample relationships. Poisson Transfer Network (PTN) [21] incorporates some relational understanding between labeled and unlabeled samples. However, it lacks the capacity to effectively mine inter- and intra-class relationships. Instance Credibility Inference (ICI) [22] ranks pseudo-labels based on credibility and sparsity, but neglects the issue of low-quality pseudo-labels. Cluster-FSL [4] employs a multi-factor clustering (MFC) strategy, which synthesizes various information sources from labeled data to produce both soft and hard pseudo-labels. It pseudo-labels selected unlabeled samples. However, we experimentally observed that the pseudo-labeling strategy employed by it led to most of the unlabeled images becoming eligible for pseudo-labeling, which would degrade the clustering process with outlier samples or incorrectly pseudo-labeled samples. Cluster-FSL’s limitations are rooted in its inadequate handling of inter-class and intra-class relationships, which restricts its ability to differentiate effectively between distinct classes and to capture subtle variations within individual classes. This leads to unlabeled samples with incorrect pseudo-labels being considered as labeled samples and consequently limiting the model performance.
In this work, we propose to mitigate these limitations through a class-variance optimized clustering (CVOC) approach and, which aligns the embedding space according to both inter- and intra-class relationships using targeted loss functions and also by using cluster separation tuner (CST). The proposed approach also incorporates semantic features to further improve the clustering process. Further, the proposed approach performs restricted pseudo-labeling and allocates pseudo-labels exclusively to unlabeled samples with low entropy predictions, reducing the possibility of noisy pseudo-labels affecting the model performance.
3 Problem Setting
In the few-shot learning (FSL) setting, there are two sets of non-overlapping classes: the base classes $C_{b}$ and the novel/few-shot classes $C_{f}$. The base classes have a sufficient number of training samples, while the novel classes have very few training samples. The dataset is split into the train set $D_{tr}$, the validation set $D_{va}$, and the test set $D_{tt}$. The train set $D_{tr}$ only contains samples from the base classes, while the validation $D_{va}$ and test $D_{tt}$ sets contain samples from the novel classes without any overlap between the classes in $D_{va}$ and $D_{tt}$. Data is provided as $N$-way $K$-shot episodes, where $N$ refers to the number of classes in the episode, and $K$ refers to the number of support/training samples from each of these classes. The support set can be denoted as $S={(x^{s}_{i},y^{s}_{i})}^{N\times K}_{i=0}$. The episode also contains $T$ randomly selected query/test samples from each of these classes on which the model can be evaluated. The query set can be denoted as $Q={(x^{q}_{i},y^{q}_{i})}^{N\times T}_{i=0}$. The episodes from the train set contain a support set $S$ and a query set $Q$ from $N$ randomly selected classes from among the classes in the train set. Similarly, the episodes from the test set contain a support set $S$ and a query set $Q$ from $N$ randomly selected classes in the test set. In this paper, we address the semi-supervised FSL setting. In this setting, apart from the above episodic data, we also have access to $u$ randomly selected samples from the classes in the episode that are not labeled. The unlabeled set can be denoted as $U={(x_{i}^{a})}^{N\times u}_{i=0}$. This leads to a setting where we have very few training samples for the classes in the episode but a good quantity of unlabeled samples.
4 Methodology
4.1 Components
The proposed approach involves training a visual feature extractor network ($M$) and a semantic injection network ($S$). The visual feature extractor $M$ is a simple deep neural network that, once trained, should be able to extract good features from images. The semantic injection network contains an encoder network ($E_{s}$) and a decoder network ($D_{s}$). Once trained, the semantic injection network $S$ should be able to map features extracted by the trained visual feature extractor from an image concatenated with the semantic feature for the corresponding class to a joint embedding space where the joint feature embedding is close to the image feature prototype of the same class.
4.2 Pre-training the Visual Feature Extractor
The visual feature extractor $M$ is first pre-trained on the train set $D_{tr}$, containing the base classes, in a regular batch-wise supervised setting. Following existing methods, we train $M$ simultaneously on the image label classification task and the rotation prediction task (see Eq. 1), since this joint training has been shown to improve the generalization power of the network. We add an image label classification head $F_{c}$ and a rotation angle classification head $F_{r}$ to $M$ for this training process.
Figure 1: Illustration of the fine-tuning phase. The support and query sets are passed through a visual feature extractor and embedding propagation. The resulting embeddings are processed in parallel by two modules: (i) Class-Variance Optimized Clustering (CVOC), which refines cluster assignments by leveraging intra- and inter-class distance regularization, followed by the Cluster Separation Tuner (CST) to enhance prototype separation before generating soft labels for query samples; and (ii) Label Propagation (LP), which also propagates support labels to query embeddings.
$L_{pt}(x_{i},y_{i},r_{i})=L_{\mathrm{ce}}\bigl{(}F_{c}\bigl{(}M(x_{i})\bigr{)},y_{i}\bigr{)}+L_{\mathrm{ce}}\bigl{(}F_{r}\bigl{(}M(x_{i})\bigr{)},r_{i}\bigr{)}\,.$ (1)
Where, $y_{i}$ and $r_{i}$ refer to the image label and rotation label for image $x_{i}$, respectively. $L_{ce}$ refers to the cross-entropy loss.
4.3 Finetuning the Visual Feature Extractor with Class-Variance Optimized Clustering
The pre-trained visual feature extractor $M$ is fine-tuned on episodes from the train set $D_{tr}$. In the few-shot regime, with only a handful of labeled support examples in each episode, aggressive augmentations greatly increase input diversity and prevent the model from simply memorizing the small support set. Following recent SSFSL methods [4], we begin by applying RandAugment [23] to every support image to guard against over-fitting. After extracting embeddings for both these augmented support samples and the query set with our visual feature extractor $M$, we refine all features using embedding propagation [3, 4], a technique commonly used in SSFSL to obtain manifold-smoothed embeddings. Finally, the propagated representations are fed into our Class-Variance Optimized Clustering (CVOC) module (Fig. 1). Then, based on the support set, the CVOC module and label propagation (LP), obtain the soft labels of the query set samples. Label propagation [24] is a graph-based technique commonly used in SSFSL to “propagate" labels from labeled samples to unlabeled samples, utilizing the similarity between the samples as weights for edges between the graph nodes.
4.3.1 Class-Variance Optimized Clustering and Cluster Separation Tuner
The proposed class-variance optimized clustering procedure utilizes an improved clustering procedure and a cluster separation tuner (CST) to get a more robust clustering of the labeled and unlabeled image features in the episode. For every class label $c\in\{1,\dots,C\}$ we build a factor dictionary
$\mathbf{F}_{c}=\Bigl{[}\underbrace{\mathbf{s}^{(c)}_{1},\dots,\mathbf{s}^{(c)}_{K}}_{\text{support}},\;\underbrace{P_{0}(c)}_{\text{initial prototype}}\Bigr{]}\in\mathbb{R}^{d\times(K+1)},$ (2)
where $\mathbf{s}^{(c)}_{k}\in\mathbb{R}^{d}$ is the $k$-th support embedding of class $c$ and $P_{0}(c)$ is their mean. Given a feature vector $x$, its representation weights (solution in the collaborative representation [25]) and reconstruction distance are
$\beta^{\!*}_{c}(x)=\arg\min_{\beta}\bigl{\|}x-\mathbf{F}_{c}\beta\bigr{\|}_{2}^{2}+\lambda\|\beta\|_{2}^{2},\quad d_{\text{rec}}(x,c)=\bigl{\|}x-\mathbf{F}_{c}\beta^{\!*}_{c}(x)\bigr{\|}_{2}^{2},$ (3)
The reconstruction-based distance captures the true shape and spread of each category by leveraging both the cluster center and representative labeled samples. This richer, multi-prototype view helps us account for within-category heterogeneity and yields more accurate, robust assignments for unlabeled data; compared to traditional clustering techniques. But $d_{\text{rec}}$ alone may be ambiguous when class subspaces overlap, so we regularise it with two Euclidean terms. Let $P(c)$ denote the current prototype of class $c$ (updated after every assignment step) and $\operatorname{ep}\!\bigl{(}M(x)\bigr{)}$ the embedding–propagated feature of $x$ as in [3]. Using $\mathcal{S}$ to denote the support set (expanded support set through augmentation), we define
$L_{\text{intra}}=\frac{1}{|\mathcal{S}|}\sum_{(x_{i},y_{i})\in\mathcal{S}}\bigl{\|}\operatorname{ep}\!\bigl{(}M(x_{i})\bigr{)}-P(y_{i})\bigr{\|}_{2}^{2},$ (4)
$L_{\text{inter}}=\frac{1}{|\mathcal{S}|}\sum_{(x_{i},y_{i})\in\mathcal{S}}\frac{1}{C-1}\sum_{\begin{subarray}{c}j=1\\ j\neq y_{i}\end{subarray}}^{C}\bigl{\|}\operatorname{ep}\!\bigl{(}M(x_{i})\bigr{)}-P(j)\bigr{\|}_{2}^{2},$ (5)
Here $C$ is the number of classes in the episode. The intra-class term $L_{\text{intra}}$ computes, for every support sample, the squared Euclidean distance between its embedding (after embedding propagation) and the prototype of its own class, and then averages these distances over the entire support set. Conversely, the inter-class term $L_{\text{inter}}$ takes each support sample, measures its squared distance to every other class prototype, averages those distances across the $C{-}1$ “foreign” classes, and finally averages the resulting values over all support samples.
The distance used both for cluster assignment and for the logits that supervise $M$ is computed as follows.
$d(x,c)\;=\;d_{\text{rec}}(x,c)\;+\;w_{\text{intra}}\,L_{\text{intra}}\;-\;w_{\text{inter}}\,L_{\text{inter}}$ (6)
with weights $w_{\text{intra}},w_{\text{inter}}$ balancing the pull–push forces. Basically, the regularizer terms shrink intra-class spread and enlarge prototype margins, leading to cleaner pseudo labels and superior few-shot accuracy relative to the previous methods (see Table 5).
Each sample $x$ (both support and query) is (re)labeled with the class that minimizes the above distance. After that, every prototype is reset to the mean of its current members. Next, we refine this set with a lightweight cluster separation tuner (CST) modelled after the continuous firefly heuristic [26]. At every CST iteration, we compute a brightness score for each prototype.
$B(c)\;=\;-w_{\text{intra}}\,L_{\text{intra}}(c)+\;w_{\text{inter}}\,L_{\text{inter}}(c)-\!\!\sum_{\begin{subarray}{c}c^{\prime}\neq c\\ \|P(c)-P(c^{\prime})\|_{2}<\varepsilon\end{subarray}}\bigl{(}\varepsilon-\|P(c)-P(c^{\prime})\|_{2}\bigr{)}^{2}$ (7)
where the first two terms reuse the same intra- and inter-class distances, but taken classwise rather than episode-wise here, and the last term penalises prototypes that sit closer than an $\varepsilon$ margin to any neighbour. If prototype $P(c)$ is dimmer than $P(c^{\prime})$ ($B(c)<B(c^{\prime})$) we move it towards the brighter neighbour with an attractiveness
$\beta=\beta_{0}\exp\!\bigl{(}-\gamma\|P(c)-P(c^{\prime})\|_{2}^{2}\bigr{)},$ (8)
yielding the update
$P(c)\;\leftarrow\;P(c)+\beta\bigl{[}P(c^{\prime})-P(c)\bigr{]}+\alpha\,\bigl{(}\text{rand}[-\tfrac{1}{2},\tfrac{1}{2}]\bigr{)}.$ (9)
Here $\beta_{0}$ sets the maximum attraction, $\gamma$ controls its spatial decay, and the Gaussian–type exploration term of amplitude $\alpha$ (decayed each loop by $0.995$) prevents premature convergence. Iterating this procedure for a small, fixed number of loops maximizes prototype separation while keeping each class center tethered to its own support cloud. For additional details, implementation specifics, and related experiments, please refer to the supplementary materials.
To predict labels for each query embedding $x_{q}$, we first compute its per-class reconstruction residuals $d_{\mathrm{rec}}(x_{q},c)$ as in Eq. (3). These distances are then converted into logits as follows.
$\ell(x_{q},c)\;=\;-\log\bigl{(}d_{\mathrm{rec}}(x_{q},c)+\varepsilon\bigr{)}\,,$ (10)
They are then normalized via softmax to obtain class probabilities as follows.
$p(c\mid x_{q})\;=\;\frac{\exp\!\bigl{(}\ell(x_{q},c)\bigr{)}}{\sum_{c^{\prime}}\exp\!\bigl{(}\ell(x_{q},c^{\prime})\bigr{)}}\,,$ (11)
Finally the predicted label is obtained as follows
$\hat{y}_{q}\;=\;\arg\max_{c}\,p(c\mid x_{q})\,.$ (12)
4.3.2 Episodic fine-tuning losses
Once CVOC has delivered a set of refined prototypes and the corresponding query logits $Q_{\text{cvoc}}\!\in\!\mathbb{R}^{(N_{q}C)\times C}$, we optimise the visual extractor $M$ with three episode-level losses:
The CVOC logits are compared with the ground-truth query labels through a cross-entropy term
$L_{\mathrm{cvoc}}=\operatorname{CE}(Q_{\text{cvoc}},y_{q}).$ (13)
In parallel, we pass the same query embeddings through the label propagation component  [24], producing logits $Q_{\text{lp}}$; the resulting loss is as follows.
$L_{\mathrm{lp}}=\operatorname{CE}(Q_{\text{lp}},y_{q}).$ (14)
To stabilize the training, we also apply the classifier head to every support and query embedding, yielding the following loss.
$L_{\mathrm{cls}}=\operatorname{CE}(Q_{\text{all}},y_{\text{all}}).$ (15)
The total objective is a convex combination as shown below.
$\mathcal{L}_{\text{episode}}=w_{\text{cls}}\,L_{\mathrm{cls}}+w_{\text{fs}}\Bigl{(}\,\eta\,L_{\mathrm{cvoc}}+(1-\eta)\,L_{\mathrm{lp}}\Bigr{)},$ (16)
where $w_{\text{cls}}$ controls the contribution of the auxiliary episode-wide term, $w_{\text{fs}}$ the overall few-shot weight, and $\eta\in[0,1]$ trades off CVOC and label propagation for the query samples. All weights are fixed on a validation split and kept constant for every episode.
4.4 Training Semantic Injection Network
Figure 2: (Left) Semantic Injection Network. (Right) Evaluation using semantic injection and restricted pseudo-labeling.
[…]
$L_{s}=L^{jep}_{s}+L^{fr}_{s}+L^{sr}_{s}$ (17)
$L^{jep}_{s}=F_{L1}\bigl{(}E_{s}(P_{c_{j}},s_{c_{j}}),P_{c_{j}}\bigr{)}$ (18)
$P_{c_{j}}=\frac{1}{N_{c_{j}}}\sum_{(x_{i},y_{i})\in D_{tr}}\mathbbm{1}(y_{i}=c_{j})\,\operatorname{ep}\bigl{(}M(x_{i})\bigr{)}$ (19)
$P^{r}_{c_{j}},\,s^{r}_{c_{j}}=D_{s}\bigl{(}E_{s}(P_{c_{j}},s_{c_{j}})\bigr{)}$ (20)
$L^{fr}_{s}=F_{L1}(P^{r}_{c_{j}},P_{c_{j}})\quad L^{sr}_{s}=F_{L1}(s^{r}_{c_{j}},s_{c_{j}})$ (21)
Where, $F_{L1}$ refers to L1-loss, $P_{c_{j}}$ refers to the class $c_{j}$ prototype, $N_{c_{j}}$ refers to the number of labeled samples belonging to class $c_{j}$. We introduce a simple, automated pipeline for extracting high-quality, visually informative class descriptions. Please refer to the supplementary material for details.
4.5 Evaluation with Semantic Injection and Restricted Pseudo-Labeling
After pre-training and episodic fine-tuning, we evaluate on $N$-way $K$-shot test episodes composed of a labeled support set $\mathcal{S}_{\text{test}}$, an unlabeled pool $\mathcal{U}_{\text{test}}$, and a query set $\mathcal{Q}_{\text{test}}$. We first run CVOC on $\mathcal{S}_{\text{test}}\!\cup\!\mathcal{U}_{\text{test}}$ to obtain visual centres $P(y)$. For each class $y$ we encode its textual description with the CLIP text encoder and obtain a semantic vector $t_{y}$ [27]. Concatenating $P(y)$ with $t_{y}$ and processing the pair through the semantic-injection encoder $E_{s}$ yields the refined prototype
$P^{\text{ref}}(y)=s\,P(y)+(1-s)\,E_{s}\bigl{(}P(y),t_{y}\bigr{)},$ (22)
where $s\in[0,1]$ is a fixed blend weight. This constitutes the overall procedure of our semantic anchor module — aligning visual prototypes via the semantic injection network and prototype correction according to Eq. (22).
For each $x_{i}^{u}\in\mathcal{U}_{\text{test}}$ we compute the entropy $H(x_{i}^{u})$ of its softmax logits and retain only the lowest-entropy $k\%$. The selected subset and its hard pseudo-labels are
$\widetilde{\mathcal{U}}_{\text{test}}=\Bigl{\{}\!\bigl{(}x_{i}^{u},\tilde{y}_{i}\bigr{)}\mid H(x_{i}^{u})\leq H_{k}\Bigr{\}},\qquad\tilde{y}_{i}=\arg\max_{c}\text{logit}_{c}(x_{i}^{u}),$ (23)
where $H_{k}$ is the entropy threshold that keeps exactly $k\%$ of $\mathcal{U}_{\text{test}}$.
The support set is then expanded with these high-confidence examples:
$\mathcal{S}^{0}=\mathcal{S}_{\text{test}}\cup\widetilde{\mathcal{U}}_{\text{test}}.$ (24)
Balanced label propagation [24] is applied to the features of $\mathcal{S}^{0}\cup\mathcal{Q}_{\text{test}}$, using the labels in $\mathcal{S}^{0}$ as anchors. The propagated soft labels for $\mathcal{Q}_{\text{test}}$ are finally hardened by selecting the highest-probability class, yielding the episode predictions (Fig. 4).
4.6 Quality of Representation Space
Prior methods in SSFSL often commit to either global prototype-based structures (e.g., ProtoNet [10], Soft K-Means [28]) or reconstruction-based alignments (e.g., MFC [4]). Prototype-based methods, while offering clear global decision boundaries, can falter in capturing fine-grained intra-class variations and local geometric nuances. Conversely, reconstruction-based clustering excels at modelling local manifolds but often mis-assign samples when two class subspaces overlap, because they ignore where class centroids sit in the broader topology. We overcome this dichotomy by fusing the reconstruction based distance with explicit Euclidean regularisers: an intra-class term that contracts each cluster and an inter-class term that repels different centroids. A point is accepted by a class only if it is simultaneously well reconstructed by that class’s subspace and lies near its prototype while far from others. This fusion retains the fine-grained shape of each class manifold while simultaneously positioning the manifolds far enough apart to maintain clear global separation. In the fine-tuning stage, the cluster separation tuner (CST) adjusts each centroid to maximize intra-class compactness and inter-class separation.
Next, we nudge each visual prototype a small step toward the CLIP [27] text embedding of its class description using our semantic injection network. This language-based anchor acts like a compass: it steers prototypes away from visually misleading regions (e.g., background clutter or occlusion) and pulls them toward positions that reflect the class’s true meaning. That is particularly valuable when we have only a handful of labeled examples.
Collectively, the geometry-aligned fusion, semantic anchoring, and margin polishing generate a latent space that is tighter within classes, farther between classes, and conceptually aligned with the class description. This enhanced discriminability yields substantially higher pseudo-label accuracy than prior approaches (see Table 5, where we compared our method against Kmeans and MFC [4]), which in turn drives significant gains in FSL performance.
Table 1: Classification accuracy (%) on the miniImageNet dataset.
ResNet-12 WRN-28-10
Method 1-shot 5-shot Method 1-shot 5-shot
TADAM 58.50$\pm$0.3076.70$\pm$0.30LEO 61.76$\pm$0.0877.59$\pm$0.12%
CAN 67.19$\pm$0.5580.64$\pm$0.35wDAE-GNN 62.96$\pm$0.6278.85$\pm$0.10%
LST 70.10$\pm$1.9078.70$\pm$0.80EPNet 80.74$\pm$0.8584.34$\pm$0.53%
EPNet 66.50$\pm$0.8981.06$\pm$0.60TransMatch 62.93$\pm$1.1182.24$\pm$0.59%
TPN 59.46% 75.65% ICI 71.41% 81.12%
PLAIN 74.38$\pm$2.0682.02$\pm$1.08EPNet-SSL 79.22$\pm$0.9288.05$\pm$0.51%
EPNet-SSL 75.36$\pm$1.0184.07$\pm$0.60PTN 81.57$\pm$0.9487.17$\pm$0.58%
cluster-FSL 77.81$\pm$0.8185.55$\pm$0.41cluster-FSL 82.63$\pm$0.7989.16$\pm$0.35%
CVOCSemRPL 84.51$\pm$0.5486.95$\pm$0.39CVOCSemRPL 85.94$\pm$0.5491.03$\pm$0.34%
Table 2: Classification accuracy (%) on the tieredImageNet dataset
ResNet-12 WRN-28-10
Method 1-shot 5-shot Method 1-shot 5-shot
MetaOpt-SVM 65.99$\pm$0.7281.56$\pm$0.53Soft K-Means 51.52$\pm$0.3670.25$\pm$0.31%
CAN 73.21$\pm$0.5884.93$\pm$0.38LEO 61.76$\pm$0.0878.43$\pm$0.12%
LST 77.70$\pm$1.6085.20$\pm$0.80wDAE-GNN 68.16$\pm$0.4682.96$\pm$0.51%
EPNet 76.53$\pm$0.8384.80$\pm$0.98EPNet 78.50$\pm$0.3188.90$\pm$0.57%
PLML-EP 79.62% 86.69% ICI 85.44% 89.12%
PLAIN 82.91$\pm$2.0989.28$\pm$1.25EPNet-SSL 83.68$\pm$0.6989.00$\pm$0.52%
EPNet-SSL 81.79$\pm$0.9788.45$\pm$0.61PTN 84.70$\pm$1.5489.14$\pm$0.73%
cluster-FSL 83.89$\pm$0.8189.94$\pm$0.46cluster-FSL 85.74$\pm$0.7690.18$\pm$0.43%
CVOCSemRPL 85.55$\pm$0.6591.13$\pm$0.31CVOCSemRPL 89.06$\pm$0.5891.53$\pm$0.32%
5 Experiments
The SSFSL methods are evaluated on the miniImageNet [29], tieredImageNet [28], and CUB [30] datasets. Please refer to the supplementary material for further details on the datasets used, the implementation details, and the compared methods.
5.1 Results
We conducted experiments under 5-way 1-shot and 5-way 5-shot settings, on the miniImageNet, tieredImageNet, and CUB-200-2011 datasets, using WRN-28-10 and ResNet-12 backbones. For the miniImageNet dataset (Table 1), our CVOCSemRPL model outperformed cluster-FSL [4], achieving a 6.7% and 1.4% improvement in 1-shot and 5-shot settings with ResNet-12, and a 3.31% and 1.87% improvement with WRN-28-10. On the tieredImageNet dataset (Table 2), CVOCSemRPL outperforms the closest method by margins of 1.66% and 1.19% for 1-shot and 5-shot cases with ResNet-12, and by a margin of 3.32% for 1-shot and 1.35% for 5-shot settings with WRN-28-10. On the CUB-200-2011 dataset (Table 3), CVOCSemRPL outperformed EPNet [3] and cluster-FSL [4] by significant margins. All experiments were conducted on 1000 few-shot tasks with 100 unlabeled data, and 95% confidence intervals are reported.
Table 3: Classification accuracy (%) on the CUB dataset.
ResNet-12 WRN-28-10
Method 1-shot 5-shot 1-shot 5-shot
EPNet 82.85$\pm$0.8191.32$\pm$0.4187.75$\pm$0.7094.03$\pm$0.33%
cluster-FSL 87.36$\pm$0.7192.17$\pm$0.3191.80$\pm$0.5895.07$\pm$0.23%
CVOCSemRPL (ours) 88.63$\pm$0.6393.37$\pm$0.1593.15$\pm$0.4196.25$\pm$0.24%
Table 4: Impact of the intra-inter class distance in finetuning and semantic injection in the testing phase on miniImageNet.
Components Accuracy
Intra–inter Semantic 5w1s 5w5s
class dist injection
$\times$ $\times$ 78.31$\pm$0.7985.75$\pm$0.41%
$\checkmark$ $\times$ 80.29$\pm$0.5485.91$\pm$0.79%
$\times$ $\checkmark$ 82.53$\pm$0.5286.45$\pm$0.41%
$\checkmark$ $\checkmark$ 84.51$\pm$0.5486.95$\pm$0.39%
Table 5: Pseudo-labeling accuracy (%) on miniImageNet.
Method 5-way 5-shot 5-way 1-shot
ResNet-12 WRN-28-10 ResNet-12 WRN-28-10
K-Means 84.13% 87.62% 76.67% 81.39%
MFC 85.04% 88.38% 77.26% 81.90%
CVOC 86.65% 90.21% 83.98% 85.41%
Figure 3: Effect of cluster separation tuner on the miniImageNet dataset 5w1s and 5w5s using ResNet12.
5.2 Ablation Experiments
We experimentally analyze the contributions of components of our method on the miniImageNet dataset and ResNet12 architecture. The results in Table 5 indicate that removing the semantic injection, or intra-inter class distance, significantly degrades the performance of our method. Fig. 3 indicates the utility of CST. Please refer to the supplementary material for other ablation experiments.
6 Conclusion
In this work, we proposed a novel SSFSL approach that performs improved clustering-based pseudo-labeling that promotes better intra/inter-class alignment and leverages semantic anchoring. Extensive experiments demonstrate that our method significantly outperforms recent state-of-the-art approaches. Limitations: Our approach assumes that both labeled and unlabeled data lie on a shared manifold, which is crucial for accurate pseudo-labeling. Our semantic anchoring technique relies heavily on the quality of raw class descriptions, which must provide rich visual characteristics to enable effective discrimination, and the choice of embedding model used to encode these descriptions. Consequently, this semantic anchor module may not work very well in complex datasets where the visual features of different classes are highly similar. This module also introduces some computational overhead.
Supplementary Material
7 Experimental Details
7.1 Datasets
The miniImageNet dataset [29] is a subset of ImageNet [31], consisting of 60,000 images, from 64 training classes, 16 validation classes, and 20 testing classes. The tieredImageNet dataset [28], is another subset of ImageNet [31] that is designed specifically for few-shot learning applications. This dataset of 779,165 images, includes a hierarchical structure with 34 superclasses divided into 608 fine-grained categories, encompassing a wide range of natural and artificial objects, animal species, and other general image classes. The dataset is partitioned into 20 training superclasses (covering 351 classes), 6 validation superclasses (97 classes), and 8 test superclasses (160 classes). The CUB dataset [30], derived from the fine-grained Caltech-UCSD Birds-200-2011 dataset [32], contains 200 classes representing 11,788 images of birds. This dataset is partitioned into 100, 50 and 50 classes for the base, validation, and novel class split.
7.2 Implementation Details
7.2.1 Pretraining
During the pre-training stage, the CVOCSemRPL model is trained for 200 epochs using a batch size of 128, utilizing all available training categories and data. A dropout rate of 0.1, weight decay of 0.0005, and momentum of 0.9 are applied. The model is optimized using the stochastic gradient descent (SGD) algorithm with an initial learning rate set to 0.1. If the validation loss does not improve for 10 consecutive epochs, the learning rate is reduced by a factor of 10. The output dimensionality of the feature representations depends on the backbone: it is 512 for ResNet-12 and 640 for WRN-28-10.
7.2.2 Finetuning
In the fine-tuning stage, the CVOCSemRPL model is trained for 200 epochs over 600 iterations. The optimization employs stochastic gradient descent (SGD) with a momentum of 0.9 and a weight decay of 0.0005. The learning rate is initially set to 0.001, and it is reduced by a factor of 10 when the validation performance plateaus. Episodic training is done at this stage.
7.2.3 Semantic Injection Network Training
The Semantic Injection Network is trained separately after episodic fine-tuning. We train for 100 epochs with a batch size of 128 support–query pairs drawn from all training classes. The input–output dimensions are 512 for image features when using ResNet-12, and 640 when using WRN-28-10 as the backbone. In both cases, the text features (semantic embeddings from CLIP)( [27]) have a dimensionality of 512. A hidden layer of size 4096 is used. Training is performed using the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$, weight decay of $1\times 10^{-4}$, and a dropout rate of 0.1. The learning rate is decayed by a factor of 0.1 every 30 epochs. The total loss is the mean of three L1 components: prototype reconstruction loss, semantic feature reconstruction loss, and joint embedding prototype loss, as detailed in the main paper. The checkpoint with the lowest validation reconstruction loss is selected for downstream few-shot evaluation.
7.2.4 Testing
During the testing phase, the CVOCSemRPL model is evaluated over 1000 episodes, and the mean classification accuracy across these episodes is reported as the final performance metric. Each few-shot task consists of 5 classes (N = 5). For every class, there are K = 1 and K = 5 support samples, along with 15 query samples (q = 15) per class. For each category, the number of unlabeled samples $u$ is 100. The temperature parameter $\tau$ is fixed at $0.1$, and the additional hyperparameter $\varepsilon$ is configured as $0.01$. The clustering process is performed over $10$ iterations.
All the hyperparameters were selected through extensive tuning across all three stages: pretraining, finetuning, and testing.
7.2.5 Compute Resources
All experiments were conducted on a single NVIDIA A6000 GPU with 48GB of VRAM. The pretraining phases utilized up to 30GB of GPU memory, while the subsequent fine-tuning and testing stages required approximately between 3GB to 8GB.
7.3 Compared Methods
Our comparative experiments include several established baselines as well as our proposed model, CVOCSemRPL. These include cluster-FSL [4], PLML+EP [1], EPNet [3], and ICI [22], along with state-of-the-art few-shot learning methods such as TADAM [11], MTL [2], MetaOpt-SVM [9], CAN [33], LST [34], and LEO [7].
We also evaluate against semi-supervised techniques like TPN [18], TransMatch [13], and PLAIN [35], as well as graph-based models such as wDAE-GNN [8]. Furthermore, comparisons are made with PTN [21] and clustering-based methods introduced by Ren et al. [28], including Soft K-Means, Soft K-Means+Cluster, and Masked Soft K-Means.
Unlike cluster-FSL, which primarily relies on quantitative evaluation of classification accuracy, our proposed method incorporates semantic context, both inter-class and intra-class variance, and unlabeled samples with low prediction entropy to enhance testing performance.
8 More details on CVOC
8.1 Algorithmic Details
8.1.1 Finetuning
In the fine-tuning stage, we invoke Alg. 1, which in each episode computes the reconstruction distance of all support and query embeddings against class-specific dictionaries, applies intra- and inter-class distance regularization and a Cluster Separation Tuner (CST, see Alg. 2) to refine the prototypes, and finally derives query logits (soft labels) from the updated prototypes. Experimentally, CST improves classification accuracy, but introduces additional computational overhead: the algorithm with CST takes approximately 45–46 seconds per epoch (training + validation) on average, whereas without CST it takes only 20–22 seconds per epoch.
Algorithm 1 Class–Variance–Optimised Clustering (CVOC) in the finetuning stage
1:labelled support embeddings $S\!=\!\{(\mathbf{s}_{i},y_{i})\}_{i=1}^{N_{s}}$, query embeddings $Q\!=\!\{\mathbf{q}_{j}\}_{j=1}^{N_{q}}$, intra/inter weights $w_{\mathrm{intra}},\,w_{\mathrm{inter}}$, ridge coefficient $\lambda$, maximum loops $T$, small constant $\varepsilon>0$, softmax temperature $\tau$
2:query logits $Q_{\text{cvoc}}\in\mathbb{R}^{N_{q}\times C}$ and hard labels $\hat{y}$
3:Initialise prototype of each class $c$ as the mean of its support: $\mu_{c}\leftarrow\frac{1}{|S_{c}|}\sum_{\mathbf{s}_{i}\in S_{c}}\mathbf{s}_{i}$
4:$A\leftarrow S\cup Q$triangleright concatenate support and query sets
5:for $t=1$ to $T$ do $\triangleright$ outer refinement loop
6:    for $c=1$ to $C$ do
7:         $\mathbf{F}_{c}\leftarrow\bigl{[}\mathbf{s}^{(c)}_{1},\dots,\mathbf{s}^{(c)}_{|S_{c}|},\;\mu_{c}\bigr{]}$
8:    end for
9:    //  Calculate Reconstruction Distance
10:    for all $\mathbf{a}\in A$ do
11:         $d_{\mathrm{rec}}(a,c)\leftarrow\text{resulting reconstruction distance}$
12:    end for
[…]
15:    //  Prototype update
16:    for all $\mathbf{a}\in A$ do
17:         $d(a,c)\leftarrow d_{\mathrm{rec}}(a,c)+w_{\mathrm{intra}}\,L_{\mathrm{intra}}-w_{\mathrm{inter}}\,L_{\mathrm{inter}}$
18:    end for
19:    Assign $A_{\mathrm{lbl}}(\mathbf{a})\leftarrow\arg\min_{c}d(a,c)$
20:    for $c=1$ to $C$ do
21:         $\mu_{c}\leftarrow\frac{1}{|A_{c}|}\sum_{\mathbf{a}\in A_{c}}\mathbf{a}$triangleright mean of current members
22:    end for
23:    $\{\mu_{c}\}\leftarrow\textsc{ClusterSeparationTuner}\bigl{(}\{\mu_{c}\}\bigr{)}$triangleright margin polishing
24:    if prototypes converge then
25:         break
26:    end if
27:end for
28://  Final logits for queries
29:for all $\mathbf{q}_{j}\in Q$ do
30:    Recalculate $d_{\mathrm{rec}}(\mathbf{q}_{j},c)$ with updated prototypes
31:    $\ell(\mathbf{q}_{j},c)\leftarrow-\log\bigl{(}d_{\mathrm{rec}}(\mathbf{q}_{j},c)+\varepsilon\bigr{)}$
32:end for
33:$Q_{\text{cvoc}}\leftarrow\mathrm{softmax}\bigl{(}\ell/\tau\bigr{)},\quad\hat{y}\leftarrow\arg\max_{c}Q_{\text{cvoc}}$
34:return $Q_{\text{cvoc}},\hat{y}$
Algorithm 2 Cluster Separation Tuner (CST)
1:initial prototypes $\{P(c)\}_{c=1}^{C}$, support embeddings $S$ with labels $y$, weights $w_{\mathrm{intra}},w_{\mathrm{inter}}$, penalty margin $\varepsilon$, parameters $\beta_{0},\gamma,\alpha$, iterations $T$
2:refined prototypes $\{P(c)\}$
3:for $t=1$ to $T$ do
4:    // Brightness evaluation
5:    for $c=1$ to $C$ do
6:         $\mathcal{S}_{c}\leftarrow\{\mathbf{s}\in S\mid y(\mathbf{s})=c\}$
7:         if $\mathcal{S}_{c}=\varnothing$ then
8:             $B_{c}\leftarrow-10^{6}$triangleright $\triangleright$ empty classes stay dim
9:         else
10:             $L_{\mathrm{intra}}(c)\;\leftarrow$ mean squared distance between $P(c)$ and samples in $\mathcal{S}_{c}$
11:             $L_{\mathrm{inter}}(c)\;\leftarrow$ mean squared distance between $\mathcal{S}_{c}$ and all other prototypes
12:             penalty $\leftarrow$ quadratic cost if any prototype lies closer than $\varepsilon$ to $P(c)$
13:             $B_{c}\leftarrow-\,w_{\mathrm{intra}}\!\cdot\!L_{\mathrm{intra}}(c)+\,w_{\mathrm{inter}}\!\cdot\!L_{\mathrm{inter}}(c)-\,$penalty
14:         end if
15:    end for
16:    // Update step
17:    for $i=1$ to $C$ do
18:         for $j=1$ to $C$ do
19:             if $B_{i}<B_{j}$ then $\triangleright$ $\triangleright$ prototype $i$ is dimmer
20:                 $\beta\leftarrow\beta_{0}\exp\!\bigl{(}-\gamma\|P(i)-P(j)\|_{2}^{2}\bigr{)}$
21:                 $P(i)\leftarrow P(i)+\beta\,[P(j)-P(i)]+\alpha\,\bigl{(}\text{rand}[-\tfrac{1}{2},\tfrac{1}{2}]\bigr{)}$
22:             end if
23:         end for
24:    end for
25:    $\alpha\leftarrow 0.995\,\alpha$triangleright $\triangleright$ anneal random exploration
26:end for
27:return $\{P(c)\}$
8.1.2 Testing
At the inference time, we first apply CVOC to the combined support and unlabeled samples to obtain a subset of high-confidence samples; these pseudo-labels are merged into the episode dictionary, and the updated combined support, unlabeled and query embeddings are passed to a Label Propagation module [24], which produces query logits and final predictions.
While semantic anchor module significantly improves accuracy, it has one drawback- computational overhead (see Table 6). For this module, we also introduce an extra hyperparameter that controls the mixing between the visual prototype and its semantic reconstruction (See Section 9.2). Apart from inference, we include a separate step to train the Semantic Injection Network, which takes approximately 30–35 seconds per epoch (training + validation) on average.
Table 6: Inference time per few-shot episode containing a total of 75 query samples on ResNet-12 (miniImageNet) measured on an NVIDIA A6000 GPU.
Method Inference Time
5w1s 5w5s
Without semantic anchor 0.181 s 0.204 s
With semantic anchor 0.330 s 0.355 s
8.2 Why Label Propagation
We experimentally demonstrate that integrating both CVOC and label propagation during the testing phase leads to a greater improvement in accuracy compared to using label propagation alone (Table 7). This is because CVOC relies heavily on labeled samples. However, the pseudo-labels obtained from the expanded support set may be noisy, which can adversely affect the performance of CVOC. Hence, employing label propagation to predict the labels of query samples during the testing phase is both a practical and effective approach.
Table 7: Comparison of classification accuracy (%) and pseudo-labeling accuracy (%) between LP (Label Propagation) and CVOC+LP methods using the ResNet-12 backbone on the miniImageNet dataset. SSFSL denotes semi-supervised few-shot learning.
Method SSFSL Acc. Pseudo Labeling Acc.
5w1s 5w5s 5w1s 5w5s
LP 75.51 84.83 63.42 80.45
CVOC + LP 84.51 86.95 83.98 86.65
8.3 Class Description Generation Pipeline
To supply the semantic-anchor module with concise yet visually informative text, we automatically build a short description for every class name [^1] WordNet glosses for each classes are automatically retrieved via the NLTK  [36] interface to WordNet [37]. by running a four–stage sequential agentic chain (see Figure 4) using the GPT-4o-mini model accessed via the OpenAI API [38]. Each stage receives the output of the previous one and is instructed with a fixed prompt; no manual intervention is required.
  1. 1.
    Writer agent — rewrites the raw WordNet definition so that it is scientifically correct, free of jargon, and $\leq 150$ words.
  2. 2.
    Filter agent — removes any behavioural, functional, or habitat information, keeping only appearance–related cues ($\leq 100$ words).
  3. 3.
    Visualizer agent — further sharpens size, shape, colour, texture, and pattern details, padding only when such cues are commonly known.
  4. 4.
    Finalizer agent — polishes the text into a single 50–70-word paragraph that provides a rich visual description of the particular class.
Writer agent prompt Please rewrite the ‘{definition}’ for ‘{class_name}’, ensuring scientific accuracy and clarity, while maintaining brevity. The definition should be detailed enough to capture essential characteristics, yet concise, within 150 words. Focus on providing clear and visually identifiable traits of the object. If the original definition includes overly technical or complex terms that aren’t needed to describe how the object looks, simplify them into plain, easy-to-understand language. Avoid adding unnecessary or unrelated information, and do not make up or guess details.
Filter agent prompt You are the filter agent. Your task is to review the following expanded definition of ‘{class_name}’. Focus on preserving only the information that directly describes the object’s visual traits and remove behaviors, habitats, functions, or technical jargon unrelated to appearance.
Definition: {rewritten_definition}
Please return the refined definition within 100 words and do not hallucinate any information.
Visualizer agent prompt You are the visualizer agent. Your task is to improve the following definition by emphasizing the key visual traits of ‘{class_name}’. Focus on attributes such as size, shape, color, texture, and patterns. If the definition is too brief or incomplete, add further relevant visual attributes based on common knowledge of the class—but do not hallucinate any information. Ensure the result is concise and highlights the most important physical characteristics.
Filtered Definition: {filtered_definition}
Please return a refined definition of no more than 100 words.
Finalizer agent prompt You are the finalizer agent. Your task is to take the following visual-focused definition of ‘{class_name}’ and polish it into a single, well-structured paragraph of 50–75 words. Make it succinct and scientifically accurate. If any details conflict, keep only the most relevant, visually descriptive information. Do not hallucinate or invent any new facts.
Visual-Focused Definition: {visual_focused_definition}
Return: A polished, 50–70-word draft-like paragraph emphasizing the key visual traits.
This fully automated pipeline yields compact, visually rich class descriptions that consistently improve semantic anchoring. We evaluated several prompt strategies - including a single-shot prompt, the unedited WordNet gloss, alternative agentic chains, and found that the four-stage pipeline described above consistently produced the strongest downstream performance.
Figure 4: Pipeline for generating visually focused class descriptions using an agentic chain powered by GPT-4o-Mini. Starting from a raw class name, a WordNet definition is retrieved and passed sequentially through four specialized agents: the Writer Agent generates a Refined Scientific Definition; the Filter Agent removes non-visual information; the Visualizer Agent enhances appearance-related cues; and the Finalizer Agent polishes the output into a concise, visually focused description. This final text is then encoded using a Text Encoder for semantic injection. Note that at each step, the class name is passed along with the definition from the previous agent.
Table 8: Semi-supervised few-shot classification on miniImageNet dataset ($5$-way 1-shot, ResNet-12 backbone). ”Num_U” represents the number of unlabeled samples.
Num_U Acc (Ours) Cluster-FSL [4]
20 80.03 72.58
50 82.81 74.83
100 84.51 77.81
200 85.30 78.31
Table 9: Semi-supervised few-shot classification on miniImageNet dataset ($5$-way 5-shot, ResNet-12 backbone).”Num_U” represents the number of unlabeled samples.
Num_U Acc (Ours) Cluster-FSL [4]
20 84.94 82.89
50 85.89 84.39
100 86.95 85.55
200 87.43 86.25
Figure 5: Accuracy comparison across different methods on the miniImageNet dataset using the ResNet-12 backbone for 5-way 1-shot and 5-way 5-shot scenarios. NN: nearest neighbour approach.
Figure 6: Impact of $s$ on 5-way 1-shot accuracy on the miniImageNet dataset using the ResNet-12 backbone
8.3.1 Effectiveness of our approach
Why begin from WordNet.
Passing only a bare class name into the writer agent can lead to confusion, because many single words have more than one meaning. For example, the word “crane” could refer to a long-legged bird or a construction machine. WordNet assigns each meaning a unique synset ID and supplies a short gloss for each sense, so we can pick the one needed for the purpose. Starting from these curated senses prevents the large language model (LLM) from mixing unrelated meanings and ensures that every class name is grounded in the correct visual category.
Why call the LLM.
Unfortunately, WordNet glosses are short (often $<\!20$ words) and rarely mention fine-grained cues such as colour patches, plumage texture, or limb shape that are crucial for few-shot recognition. We therefore prompt GPT-4o-mini to expand each gloss, injecting the missing visual details.
Why an agentic refinement chain.
Iterative self-refinement has been shown to improve factuality and task performance in LLMs [39]. So, our four-stage pipeline writer $\rightarrow$ Filter $\rightarrow$ Visualizer $\rightarrow$ Finalizer realises this idea: the Filter runs at low temperature to deterministically prune non-visual content, while the Visualizer is sampled with a slightly higher temperature to fill in plausible but commonly known traits; the Finalizer then consolidates the description into 50–70 highly informative words. Controlling generation temperature is a well-studied way of trading off creativity and determinism in LLMs.
Removing noise without losing signal.
A small amount of non-visual info can help steer the text embedding away from confusion. Aggressive pruning eliminates habitat or behavioural clauses that could hurt visual grounding (e.g. “feeds on carrion”) yet keeps concise non-visual hints that implicitly map to appearance: nocturnal $\rightarrow$ large eyes, aquatic $\rightarrow$ streamlined body, raptor $\rightarrow$ hooked beak. Such implicit cues have proved beneficial when text embeddings are later aligned with vision features .
Disambiguating look-alike classes.
Some birds look almost the same in pictures - take an ’American Crow’ and ’Common Raven’. Both are shiny black corvids, but ravens are noticeably bigger and have a thicker bill. Standard WordNet definitions do not point out these visual differences, but our process keeps them. That way, our approach learns to differentiate between the two, rather than mixing them up.
This iterative, refinement-driven expansion and denoising procedure yields semantically rich yet compact class texts that translate into stronger visual anchoring and ultimately, higher SSFSL (Semi-Supervised Few-Shot Learning) accuracy.
8.4 Effect of unlabeled samples
In this section, we analyze how our results vary with the availability of unlabeled samples. As we increase the number of unlabeled samples, we observe a corresponding improvement in performance, which is intuitive (see Tables 8 and 9). We also compare our method with Cluster-FSL [4], which represents the previous state-of-the-art architecture in semi-supervised few-shot learning prior to our work.
9 Ablation Experiments
9.1 Ablation on the contributions of the Cluster Separation Tuner and Semantic Correction
In this section, in order to isolate the contributions of the Cluster Separation Tuner (CST) and semantic correction, we disable the reconstruction-distance term and plug CST + semantic anchor directly into standard clustering methods. Here, we compare against the latest methods under the 5-way 1-shot and 5-shot tasks. Although we had to retune a few hyper-parameters - chiefly $s$ in the semantic anchor module and the number of CST iterations to fit this plug-in setup, the adjusted settings still yield competitive performance (Figure 5).
9.2 Ablation on hyperparameter $s$ in Semantic anchor
In the semantic anchor module, we introduce an extra hyperparameter ($s\in[0,1]$; as discussed in the main paper) that controls the strength of shifting visual prototypes toward semantically plausible regions, helping to alleviate issues such as occlusions. Experimentally, this module proves especially beneficial in the 5-way 1-shot setting, likely because a single support example is more susceptible to errors when that image is occluded or otherwise ambiguous.As shown in Figure 6, accuracy steadily increases with $s$, peaks at $s=0.9$, and then declines at $s=1$.
9.3 Ablation on Restricted Pseudo Labeling
At test time, we limit the augmentation of the support set by selecting only those unlabeled samples with the most confident pseudo-labels. Since pseudo-labels can be noisy, we sort unlabeled samples by their prediction entropy and pick the lowest-entropy examples, offering a simple, efficient strategy with minimal test-time overhead (see Figure 7). We observe that retaining 80% of the unlabeled samples yields the highest gains, highlighting the impact of confident pseudo-label selection.
Figure 7: Accuracy on miniImageNet as a function of RPL (%) - the proportion of unlabeled samples retained based on lowest-entropy pseudo-labels. The 5w1s and 5w5s curves correspond to 5-way 1-shot and 5-way 5-shot settings, respectively, using a ResNet-12 backbone.
Figure 8: Accuracy comparison between BERT-base and CLIP text encoders on miniImageNet in 5-way 1-shot and 5-way 5-shot settings using ResNet-12 backbone. CLIP’s multimodal pretraining yields consistently higher few-shot accuracy.
9.4 Choice of embedding model
We evaluate two text encoders for generating class-description embeddings: a vanilla BERT-base model [40] and the CLIP text encoder [27]. In both 5-way 1-shot and 5-way-5-shot settings, CLIP’s text features outperform BERT significantly (see Figure 8).
Potential reasons might be that CLIP’s text encoder is trained jointly with its image encoder on millions of image–caption pairs via a contrastive objective, grounding its representations in visual concepts. BERT, by contrast, learns only linguistic co-occurrence patterns and lacks alignment with vision. Consequently, CLIP’s text vectors naturally becomes far more effective option for bridging semantic gaps in few-shot vision tasks than a standalone language model.